# supSupplementary References

### v1

@@ 7 Conclusion
We have presented RoBBR, a benchmark for assessing the ability to judge risk of bias in biomedical papers. The benchmark is based on gold-standard risk of bias assessments from Cochrane, and uses a human-validated annotation pipeline for scalably constructing task examples. The four benchmark tasks evaluate different skills required for these assessments, including retrieval and reasoning.
Performance of state-of-the-art language models on the benchmark tasks reveals that current models fall significantly short of expert-level performance. By providing a standardized tool for measuring risk of bias judgments, the RoBBR benchmark can help guide the development of AI systems that aim to automatically assess study quality and synthesize scientific evidence. As systems that review the scientific literature are developed, it is important that they can reliably judge the strength of evidence in order to draw sound conclusions. ML systems which perform well on this benchmark could potentially reduce the time required for risk of bias assessments in systematic reviews, which currently averages nearly six hours per study Crocker et al. [2023]. We hope the RoBBR benchmark will spur further research into developing trustworthy systems for this task.
8 Limitations
The RoBBR benchmark has several limitations. The source of all bias judgments in the benchmark is Cochrane meta-analyses. While these are considered very high quality meta-analyses among biomedical researchers, the bias judgments that they contain are likely to have occasional errors. In addition, there can be disagreements in judgment among different meta-analysis authors.
The Cochrane risk of bias framework does not necessarily identify all types of bias in research studies. The framework uses a predefined set of bias categories, which may not be appropriate for all studies.
The benchmark only includes interventional studies, and therefore excludes observational studies. Assessment of risk of bias in observational studies requires a different procedure than for interventional studies. Our benchmark does not measure performance for observational studies.
Only a limited number of prompting strategies were evaluated for the LLMs. It is likely that improved prompting and task presentation would improve performance on the tasks.
Appendix: Table of Contents
  1. A.
    DatasetA
    1. 1.
      Dataset License and Code LicenseA.1
    2. 2.
      Dataset Hosting, Accessibility and MaintenanceA.2
    3. 3.
      Dataset StatisticsA.3
    4. 4.
      Dataset Collection and Processing A.4
    5. 5.
      RoBBR Structure A.5
  2. B.
    Annotation GuidelineB
  3. C.
    Task 2 Bias Retrieval Prompt and OptimizationC
    • Aspect Decomposition Prompt C.1
    • Aspect Filtering Prompt C.2
    • Aspect-to-Sentence Mapping Prompt and Optimization C.3
    • A Motivating Example C.4
  4. D.
    Task 3 Synthetic Option Generation Procedure and PromptD
  5. E.
    Experiment DetailsE
    • Prompt Optimization E.1
    • Prompts for Evaluation E.2
      • Prompt to Evaluate Study Inclusion/Exclusion (Task 1) E.2.1
      • Prompt to Evaluate Bias Retrieval (Task 2) E.2.2
      • Prompt to Evaluate Support Judgment Selection (Task 3) E.2.3
      • Prompt to Evaluate Risk Level Determination (Task 4) E.2.4
    • Results with Standard Error E.3
  6. F.
    Full Example F
    • Search Protocol F.1
    • Objective of the Meta-Analysis F.2
    • Full Paper F.3
    • Bias Name and Definition F.4
    • PICO F.5
@@ A.1 Dataset License and Code License
The RoBBR dataset is made available under CC-BY-NC. A copy of the full license can be found at https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md.
The code used in this paper is released under the MIT License. The MIT License is a permissive open-source license that allows for the free use, modification, and distribution of the code, as long as the original license is included with any derivative work. A copy of the full license can be found at https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md.
@@ A.2 Dataset Hosting, Accessibility and Maintenance
The RoBBR dataset with its meta-data is released and can be accessed freely at (https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md). We commit to regularly maintain the dataset and codebase by incorporating user feedback. We will potentially introduce more features as part of future work in the next version of RoBBR. We confirm that the current version of RoBBR will always remain accessible at the same link.
@@ A.3 Dataset Statistics
See table 6 7 8 9 for detailed test set statistics of the four tasks.
Table 6: Test set statistics for Task 1 (Study Inclusion/Exclusion)
# of Tokens in Query Correct Option Count
Dataset Points min avg max Included Excluded
Test Set 309 2685 9380 28464 128 181
Table 7: Test set statistics for Task 2 (Risk of Bias: Sentence Retrieval)
# of Tokens in the Query Sentences Covered Aspects Optimal # of sentences
Dataset Points min avg max min avg max min avg max min avg max
Test Set 513 4104 11220 21494 87 220.1 400 1 1.9 10 1 1.4 9
Table 8: Test set statistics Task 3 (Risk of Bias: Support Judgement Selection)
# of Tokens in the Query
Dataset Points min avg max
Test Set 479 3515 10583 21187
Table 9: Test set statistics for Task 4 (Risk of Bias: Risk Level Determination)
# of Tokens in Query Correct Option Count
Dataset Points min avg max Low Unclear High
Test Set 1052 3256 9763 19834 676 213 163
Table 10: Test set category statistics for Task 2, 3, and 4
Bias Type
Task Points Selection Performance Detection Reporting Attrition Other Analysis Paper specific
Task 2 513 202 81 82 35 95 23 13 7
Task 3 479 130 87 82 80 98 0 8 6
Task 4 1052 332 170 168 119 148 102 18 26
@@ A.4 Dataset Collection and Processing
[…]
@@ A.5 RoBBR Structure
  • task1_SIE_test.json and task1_SIE_dev.json: Test and dev set of task 1 Study Inclusion/Exclusion
    • paper_doi: The DOI of the paper.
    • objective: The meta-analysis objective.
    • search_protocol: search protocol information of the meta-analysis.
    • full_paper: The full paper content.
    • label: One of [included, excluded], showing whether the paper is included or excluded in the meta-analysis.
  • task2_ROBSR_test.json and task2_ROBSR_dev.json: Test and dev set of task 2 Risk of Bias Sentence Retrieval
@@
    • bias_retrieval_at_optimal_evaluation: This is a dictionary that contains the necessary information for evaluating the model’s performance on the task Bias Retrieval @Optimal.
@@
    • bias_retrieval_at_3_evaluation: This is a dictionary that contains the necessary information for evaluating your model’s performance on the task Bias Retrieval @3.
        one_selection_of_sentences: a list of 3 sentence indices. The list of sentences cover the largest number of aspects that can be covered under the restriction of 3 sentences. Note, there are potentially other lists of sentences that has size 3 and cover the same number of aspects.
        covered_aspects: the list of aspects that are covered. In this case, this list of aspects may not be all the aspects. Since in the paper, we calculate aspect recall by diving the number of aspects covered by model’s retrieved sentences against the total number of aspects, for Bias Retrieval @3, the maximum possible performance is not 100%.
  • task3_SJS_test.json and task3_SJS_dev.json: Test and dev set of task 3 Support Judgment Selection
@@
  • task4_RLD_test.json and task4_RLD_dev.json: Test and dev set of task 4 Risk Level Determination
@@ Appendix B Annotation Guideline
Appendix C Task 2 Bias Retrieval Prompt and Optimization
C.1 Aspect Decomposition Prompt
@@
C.2 Aspect Filtering Prompt
See Figure 7 for prompt to filter negations.
See Figure 8 for prompt to filter non-specific commentaries.
Figure 7: Prompt for extracting negations.
Figure 8: Prompt for filtering non-specific commentaries.
C.3 Aspect-to-Sentence Mapping Prompt and Optimization
To enhance the agreement between human annotators and the GPT4-0125-turbo model, we developed and tested various prompt engineering strategies on a development set consisting of 30 unique (aspect, paper) pairs. It is important to note that no paper or aspect from the development set overlapped with the 50 tasks in the hypothesis testing set.
GPT4-0125-preview was chosen due to its robust instruction-following capabilities and our familiarity with its performance across different settings. Our experimentation revealed that using in-context learning examples did not improve agreement rates and tended to cause overfitting. Instead, we found that embedding the same instruction at both the beginning and the end of the prompt effectively helped the model maintain focus on the task of identifying text elements that significantly cover the details specified in an aspect.
We also implemented chain-of-thought reasoning, prompting GPT4 to articulate its thought process following a specific keyword. This approach not only enhanced the quality of the model’s reasoning but also stabilized its performance and reduced common-sense errors.
To address the issue of entity forgetting in long-context tasks (a typical paper might contain around 5000 tokens), we employed a sliding window technique. Each window, containing 10 text elements with a 5-element overlap, allowed GPT4 to process and evaluate each text element within a manageable context size. The hyperparameters, window length and overlap size, were optimized using the development set. This overlapping approach ensures that each text element is evaluated twice, significantly reducing the likelihood of false negatives.
Figure 9 illustrates the final prompt template used to match text elements (such as sentences, tables, figure captions, etc.) with the aspects. Each template inputs one aspect, the 10 text elements within a sliding window, and the contextual background (specifically, the summary of the IARC paper from which the aspect was derived).
@@ C.3 Aspect-to-Sentence Mapping Prompt and Optimization
Figure 9: Prompt for generating synthetic options.
C.4 A Motivating Example
Here we provide an example of how we build the bias retrieval task.
@@ C.4 A Motivating Example
Mapping Aspects to Sentences in the paper
Utilizing the procedure described in section C.3, we map the remaining 6 aspects to all sentences in the paper Werneck et al. [2014].
We only show sentences from the paper that are matched with at least one aspect.
@@
Bias Recall @ Optimal
@@
One of our evaluation metric, Bias Recall @ Optimal, measures the the optimal number of sentences to cover all aspects. In this example, we only need two sentences, sentence 66 and sentence 73, to cover all aspects.
Appendix D Task 3 Synthetic Option Generation Procedure and Prompt

### v2

@@ 7 Conclusion
We present RoBBR, a benchmark for measuring models’ ability to assess risk-of-bias in biomedical studies. RoBBR includes two novel subtasks that evaluate a model’s retrieval and reasoning abilities. The support sentence retrieval subtask is created by our fully automatic and human-validated annotation pipeline for aligning support judgments to sentences in biomedical studies. Our analysis reveals the importance of retrieval and reasoning abilities and demonstrate their impact on the a model’s ability to assess risk-of-bias.
Limitations
While in this work all systematic reviews and biomedical studies are open-sourced (Creative Commons License or Public Domain), many other systematic reviews have licenses which restrict distributing their contents. Therefore, we share the entire codebase for creating the RoBBR benchmark and encourage researchers with access to non-open-source contents to create other versions of RoBBR.
In our evaluation of LLMs, while we have developed our prompts using RoBBR’s train sets and investigated popular techniques such as chain-of-though reasoning etc, we acknowledge the possibility that better prompting techniques could lead to slight model improvements on RoBBR.
Ethics Statement
All data in RoBBR come from open-sourced systematic reviews and biomedical studies covering the domain of evidence-based biomedicine. The authors make sure no personal information is included in RoBBR by manual inspection of all datapoints.
The RoBBR benchmark, which promotes the automation of risk-of-bias determination, might introduce over-reliance on commercial AI systems to evaluate a biomedical study’s risk-of-bias.
The benchmark only includes systematic reviews and biomedical studies in English. Systems trained and evaluated on RoBBR might disadvantage the risk-of-bias determinations for non-English biomedical studies.
@@ A.1 Dataset License and Code License
The RoBBR dataset is made available under CC-BY-NC. A copy of the full license can be found at RoBBR License.
The code used in this paper is released under the MIT License. The MIT License is a permissive open-source license that allows for the free use, modification, and distribution of the code, as long as the original license is included with any derivative work. A copy of the full license can be found at RoBBR License.
@@ A.2 Dataset Hosting, Accessibility and Maintenance
The RoBBR dataset with its meta-data is released and can be accessed freely at RoBBR License. We commit to regularly maintain the dataset and codebase by incorporating user feedback. We will potentially introduce more features as part of future work in the next version of RoBBR. We confirm that the current version of RoBBR will always remain accessible at the same link.
@@ A.3 Dataset Statistics
See Tables 9, 10, 11 for detailed statistics of the RoBBR tasks.
Statistic min avg max
Tokens per report 2,908 9,215 19,080
Sentences 87 218.6 400
Covered Aspects 1 1.9 7
Optimal # of sentences 1 1.4 5
Table 9: Statistics for the Cochrane Test Set SSR (313 data points).
Bias Type Cochrane Test SSR Cochrane Test SJS Cochr. + Non-Cochr. Test Main
Selection 157 130 933
Attrition 46 98 629
Performance 54 87 309
Detection 61 82 645
Reporting 14 80 594
Deviation 0 0 331
Table 10: Test set category statistics
Statistic Cochrane Train Non-Cochr. Test Cochrane Test
Points 774 2 489 906
Tokens per report:
min 1,991 1,488 2,274
avg 9,907 7,791 8,946
max 18,894 26,154 19,080
Label distribution:
Low 387 1,465 562
Unclear 275 574 195
High 112 450 149
Table 11: Test-set statistics for Risk-of-Bias Determination. Unclear/Some concerns for Non-Cochrane test
@@ A.4 Dataset Collection and Processing
We use BioC Comeau et al. (2019) API to download meta-reviews and papers in PMC database. We manually download the rest of the papers. We use GROBID Lopez (2008–2024) to parse papers from PDF to XML format. We use Stanza Qi et al. (2020) to split paragraphs (including table and figure captions) into sentences.
@@ A.5 RoBBR Structure
  • SSR_test.json and SSR_dev.json: Test and dev set of Support Sentence Retrieval
@@
    • bias_retrieval_at_optimal_evaluation: This is a dictionary that contains the necessary information for evaluating the model’s performance on the task Support Sentence Retrieval @ Optimal.
@@
  • SJS_test.json and SJS_dev.json: Test and dev set of Support Judgment Selection
@@
  • Main_task_test.json and Main_task_dev.json: Test and dev set of the main task
@@ Appendix B Annotation Guideline
    End of Annotation Guideline.
    B.1 Annotator Training
    The annotation teams each comprised two graduate students. To ensure a thorough understanding of bias names and risk levels, all annotators were trained and briefed on their definitions prior to beginning annotation. A key part of their preparation involved an in-depth reading of five to ten full-length meta-analyses from Cochrane Train (Main), including their extensive supplementary materials (often over 100 pages), to gain familiarity with relevant terminology and details. Although our annotation guidelines allowed for the use of Large Language Models (LLMs) to facilitate understanding of unfamiliar subject matter, annotators infrequently consulted these tools. Their preference was to use search engines (with AI overview features disabled). On the rare occasions LLMs were employed, their function was strictly confined to assisting with the comprehension of unfamiliar technical details. Annotators were monitored and were expressly forbidden from soliciting LLM decisions or opinions regarding annotation choices. It is important to note that LLMs were infrequently utilized and only as a supplementary tool for assisting the understanding of unfamiliar subject matters.
Appendix C Task Descriptions and Visualizations
We provide visual illustrations of the three tasks.
C.1 Main Task: Risk-of-Bias Determination
The model is given the whole biomedical paper, study characteristics (i.e. participants, intervention methods, comparators, outcome, etc.), the objective/topic of the systematic review, one bias name and definition, and general guideline that explains how to classify the risk level as high, low or unclear/some concern, and is asked to determine the risk level of bias. Note, the definition is based on the specific bias name and risk level. Our categorization of different bias names into six main categories does not affect the definition of individual biases. See Table 12 for examples of bias categorization.
Bias Name Bias Category
random sequence generation (selection bias) selection bias
bias arising from the randomization process selection bias
bias due to deviations from intended intervention deviation bias
bias due to missing outcome data attrition bias
incomplete outcome data (attrition bias) all outcomes attrition bias
similarity of other baseline characteristics (selection and performance bias) selection bias / performance bias
bias in measurement of the outcome detection bias
blinding of outcome assessment (detection bias) all outcomes detection bias
blinding for adverse events (performance and detection bias) performance bias / detection bias
blinding of participants and personnel (performance bias) performance bias
selective reporting (reporting bias) reporting bias
bias in selection of the reported result reporting bias
blinding (performance bias and detection bias) all outcomes performance bias / detection bias
Table 12: Examples of Bias Categorization
C.2 Subtask: Support Sentence Retrieval
When judging the risk of bias for a specific study, the review authors provide support judgments which justify their risk of bias rating. These support judgments are grounded in specific aspects of the study, which the support judgment describes. This aims to test a model’s ability to retrieve the correct source for the support judgment from a paper’s text. Note the model is instructed to retrieve no more than the optimal number of sentences. The optimal number can be calculated from Aspect Sentence Mapping and is already included in our dataset. See Figure 5 for an illustration.
C.3 Subtask: Support Judgment Selection
Figure 5: Subtask: Support Sentence Retrieval. In this task, the goal is to retrieve no more than the optimal number of sentences from the biomedical report which support a risk of bias judgment. Performance is evaluated against a support judgment, which has been split into aspects.
Figure 6: Subtask: Support Judgment Selection. In this task, the model is shown a paper, and asked which of 7 support judgments provides the best explanation of the paper’s risk of bias. Only four options are shown here for illustration purposes. Both options C and D refer to the same information from the paper, in this case the operational challenges affecting the trial’s methodology, but describe different reasoning given this information. This demonstrates that retrieving the correct information from the paper is not sufficient for solving this task.
Appendix D Support Sentence Retrieval Prompt, Optimization and Example
D.1 Aspect Decomposition Prompt
@@
D.2 Aspect Filtering Prompt
See Figure 8 for prompt to filter non-specific commentaries.
@@ D.2 Aspect Filtering Prompt
D.3 Aspect-to-Sentence Mapping Prompt and Optimization
See Figure 9 for the prompt optimized on the development set.
To enhance the agreement between human annotators and the GPT-4-0125-preview model, we developed and tested various prompt engineering strategies on a development set consisting of 30 unique (aspect, paper) pairs. It is important to note that no paper or aspect from the development set overlapped with the 50 tasks in the hypothesis testing set.
GPT-4-0125-preview was chosen due to its robust instruction-following capabilities and our familiarity with its performance across different settings. Our experimentation revealed that using in-context learning examples did not improve agreement rates and tended to cause overfitting. Instead, we found that embedding the same instruction at both the beginning and the end of the prompt effectively helped the model maintain focus on the task of identifying text elements that significantly cover the details specified in an aspect.
We also implemented chain-of-thought reasoning, prompting GPT-4 to articulate its thought process following a specific keyword. This approach not only enhanced the quality of the model’s reasoning but also stabilized its performance and reduced common-sense errors.
To address the issue of entity forgetting in long-context tasks (a typical paper might contain around 5000 tokens), we employed a sliding window technique. Each window, containing 10 text elements with a 5-element overlap, allowed GPT-4 to process and evaluate each text element within a manageable context size. The hyperparameters, window length and overlap size, were optimized using the development set. This overlapping approach ensures that each text element is evaluated twice, significantly reducing the likelihood of false negatives.
Figure 9 illustrates the final prompt template used to match text elements (such as sentences, tables, figure captions, etc.) with the aspects. Each template inputs one aspect, the 10 text elements within a sliding window, and the contextual background.
@@ D.3 Aspect-to-Sentence Mapping Prompt and Optimization
D.4 A Motivating Example
Here we provide an example of how we build the support sentence retrieval task.
@@ D.4 A Motivating Example
Mapping Aspects to Sentences in Report
Utilizing the procedure described in section D.3, we map the remaining 6 aspects to all sentences in the report Werneck et al. (2014).
We only show sentences from the report that are matched with at least one aspect.
@@
Aspect Recall Ratio @ Optimal
@@
One of our evaluation metric, Aspect Recall Ratio @ Optimal, measures the the optimal number of sentences to cover all aspects. In this example, we only need two sentences, sentence 66 and sentence 73, to cover all aspects.
Figure 8: Prompt for filtering non-specific commentaries.
Figure 9: Prompt for generating synthetic options.
Appendix E Task-specific Fine-tuning
We fine-tune Llama-3-8B on the three train sets of RoBBR’s three tasks using LoRA Hu et al. (2021), on 1 node of 8 NVIDIA H-100. Our fine-tuning hyperparameters are standard: LoRA rank =8=8, LoRA alpha =16=16, a batch size =16=16, with the AdamW optimizer (weight decay =0.01=0.01), and a cosine learning rate scheduler with 10% warmup. Llama-3-8B is fine-tuned for 1 epoch using cross entropy loss. The model is asked to predict the label (i.e., risk-level for main task, sentence indices for SSR, correct support judgment for SJS). We compared different fine-tuning techniques, and found when we train model to predict both the label and a paragraph of reasoning text, the optimal training loss is achieved on the train sets, since it increases the amount of training tokens directly go through cross entropy loss Hsieh et al. (2023). We use support judgment as proxy for reasoning text.
Appendix F Analysis: Two Types of Long-Context Problems
Figure 10: Local-connection problems are easier for LLMs.
An active research direction in long-context modeling is to categorize different types of long-context problems Li et al. (2024). Using our annotation pipeline’s Aspect-to-Sentence mapping, we can naturally categorize RoBBR’s datapoints into two types by the optimal number of support sentences used to cover all aspects.
Local-Connection Problem:
If one support sentence can cover all aspects that form the support judgment, it means this support sentence is all a model needs to reach the same support judgment. While the main task for this datapoint is still a long-context problem because RoBBR’s average document length exceeds 8k tokens, it can be solved by locating one specific sentence from the report. There are 213 local-connection datapoints in RoBBR.
Global-Connection Problem:
If the support judgment cannot be traced back to one single sentence, some degree of global understanding of the report is required. There are 693 global-connection datapoints in RoBBR.
In Figure 10, we calculate the average LLM accuracy per type, and found evidence that suggests local-connection problems are easier for both GPT-4o and Sonnet-3.5. The intuition is that a problem becomes harder when a more global understanding of the entire long context is required.
Appendix G Error Analysis
To better illustrate that GPT-4o and Sonnet-3.5 possess different reasoning strengths, we compare their reasoning traces when both are given the same ground-truth retrieval results.
Bias: Allocation concealment (selection bias)
Bias Definition and Selection Criteria for Risk Level:
Selection bias (biased allocation to interventions) due to inadequate concealment of allocations prior to assignment.
low risk: Participants and investigators enrolling participants could not foresee assignment because one of the following, or an equivalent method, was used to conceal allocation: Central allocation (including telephone, web-based and pharmacy-controlled randomization); Sequentially numbered drug containers of identical appearance; Sequentially numbered, opaque, sealed envelopes.
high risk: Participants or investigators enrolling participants could possibly foresee assignments and thus introduce selection bias, such as allocation based on: Using an open random allocation schedule (e.g. a list of random numbers); Assignment envelopes were used without appropriate safeguards (e.g. if envelopes were unsealed or non-opaque or not sequentially numbered); Alternation or rotation; Date of birth; Case record number; Any other explicitly unconcealed procedure.
unclear risk: Insufficient information to permit judgement of ’Low risk’ or ’High risk’. This is usually the case if the method of concealment is not described or not described in sufficient detail to allow a definite judgement – for example if the use of assignment envelopes is described, but it remains unclear whether envelopes were sequentially numbered, opaque and sealed.
Retrieved Sentence:
Patients were randomly assigned to the control or experimental group following a simple randomization procedure (computerized random numbers) achieved using opaque envelopes.
Physical therapy was started under consultation with the rehabilitation medicine department after wound healing, and the allocation information to each group was not provided to reduce bias.
Sonnet-3.5:
REASON: The study describes using a computerized random number generator for randomization, which is a good start for reducing selection bias. They also mention using opaque envelopes, which aligns with one of the low-risk criteria (sequentially numbered, opaque, sealed envelopes). However, the description doesn’t explicitly state that the envelopes were sequentially numbered or sealed. Additionally, while they mention that allocation information was not provided to reduce bias, it’s not entirely clear if this refers to blinding after assignment or concealment before assignment.
The use of opaque envelopes suggests an attempt at concealment, but without the full details of how these envelopes were used (sequentially numbered, sealed), we can’t be certain that the concealment was adequate to prevent foresight of assignments. The computerized randomization is a positive factor, but the implementation of the envelope system is crucial for proper concealment.
Given these factors, there is some evidence of attempts to conceal allocation, but not enough information to confidently classify it as low risk. At the same time, there’s no clear indication of high risk practices.
DECISION: unclear
GPT-4o:
REASON: The study describes the use of opaque envelopes for allocation concealment, which is a method that can prevent participants and investigators from foreseeing assignments, provided that the envelopes are sequentially numbered, opaque, and sealed. The highlighted sentences indicate that opaque envelopes were used, but do not specify whether they were sequentially numbered and sealed. However, the mention of using opaque envelopes suggests an attempt to conceal allocation, and the additional information that allocation information was not provided to reduce bias further supports the intention to prevent foreseeing assignments. Despite the lack of explicit detail about sequential numbering and sealing, the description leans towards a method that is generally considered adequate for allocation concealment.
DECISION: low
The provided example demonstrates that Sonnet-3.5 is better in reasoning compared to GPT-4o. Sonnet-3.5 correctly identifies critical information gaps in the allocation concealment process—while opaque envelopes were used, there’s no confirmation they were sealed and sequentially numbered, which are important for allocation concealment according to the decision criteria. Sonnet-3.5 also recognizes that the mention of withheld allocation information could indicate post-assignment blinding rather than proper pre-assignment concealment. In contrast, GPT-4o fails to identify these crucial gaps and makes unwarranted assumptions about concealment adequacy without supporting evidence for key criteria. This example clearly illustrates Sonnet-3.5’s better reasoning capabilities.
Appendix H SJS Synthetic Option Generation Procedure and Prompt